### v3

# One-Step Offline Distillation of Diffusion-based Models via Koopman Modeling

Abstract
Diffusion-based generative models have demonstrated exceptional performance, yet their iterative sampling procedures remain computationally expensive. A prominent strategy to mitigate this cost is distillation, with offline distillation offering particular advantages in terms of efficiency, modularity, and flexibility. In this work, we identify two key observations that motivate a principled distillation framework: (1) while diffusion models have been viewed through the lens of dynamical systems theory, powerful and underexplored tools can be further leveraged; and (2) diffusion models inherently impose structured, semantically coherent trajectories in latent space. Building on these observations, we introduce the Koopman Distillation Model (KDM), a novel offline distillation approach grounded in Koopman theory—a classical framework for representing nonlinear dynamics linearly in a transformed space. KDM encodes noisy inputs into an embedded space where a learned linear operator propagates them forward, followed by a decoder that reconstructs clean samples. This enables single-step generation while preserving semantic fidelity. We provide theoretical justification for our approach: (1) under mild assumptions, the learned diffusion dynamics admit a finite-dimensional Koopman representation; and (2) proximity in the Koopman latent space correlates with semantic similarity in the generated outputs, allowing for effective trajectory alignment. KDM achieves highly competitive performance across standard offline distillation benchmarks.
1 Introduction
Diffusion-based generative models sohl2015deep , including score-based and flow-matching variants ho2020denoising ; lipman2023flow , have become ubiquitous across a wide range of domains—from images and videos to audio and time series dhariwal2021diffusion ; ho2022video ; kong2020diffwave ; coletta2023constrained ; naiman2024utilizing ; fadlon025diffusion ; gonen2025time . These models now surpass approaches such as GANs and VAEs in terms of sample quality, while also exhibiting greater training stability. Despite these advantages, one of their key limitations remains the high computational cost associated with sampling nichol2021improved ; salimans2022progressive ; luo2023latent . Generating a single high-fidelity sample typically requires executing a lengthy iterative process, progressively refining random noise through dozens or even hundreds of model evaluations ho2020denoising .
A widely adopted strategy to address this involves minimizing the number of inference steps song2020denoising , for example by leveraging improved numerical solvers or designing more expressive noise schedules, enabling high-quality generation with significantly fewer denoising iterations. Another increasingly popular alternative is distillation salimans2022progressive , where a student model learns to emulate the behavior of a teacher model. Distillation approaches vary in supervision and setup: in online distillation, the student directly learns from the teacher’s predictions during training, with the teacher providing trajectory information across time steps. In contrast, offline distillation relies on precomputed noise–image pairs or teacher-generated samples only, where the student learns without any further access to the teacher model, including its weights, evaluations, or internal states.
Offline methods inherently offer key advantages over online tools. Training on precomputed, teacher-generated noise–image pairs avoids multi-stage training and repeated teacher queries, reducing both neural function evaluations and memory footprint geng2023one . It also enables pre-filtering or anonymization of potentially privacy violations fernandez2023privacy . Consequently, once the dataset is produced, the teacher can be discarded, shrinking the attack surface and supporting secure collaboration (e.g., outsourcing training without exposing proprietary information) (carlini2023extracting, ; hu2023membership, ; nasr2023scalable, ). Beyond privacy and efficiency, offline distillation is model-agnostic and architecture-general (as evidenced in our experiments), and it scales flexibly with parameter budgets. On the limitation side, offline setups require upfront generation and storage costs for the precomputed corpus and, lacking teacher access beyond the noise–image pairs, may constrain model expressivity and adaptability. By contrast, online methods may warm-start from teacher weights or query the teacher for stepwise supervision. Despite these trade-offs, offline distillation remains a practical, private, and sustainable approach to generative-model distillation. The two paradigms can be considered complementary, and the optimal choice depends on factors such as privacy requirements, computational cost, and deployment constraints.
Towards developing a distillation method, we identify two key observations that guide our approach: (1) While denoising-based methods leverage dynamical systems perspectives for both general generative modeling song2021score ; lipman2023flow and distillation zhang2023fast ; chen2023geometric , a wide range of powerful tools from dynamical systems theory remain underexplored; (2) Trained diffusion models impose a structured organization on the latent space, coherently mapping noisy inputs toward clean data. Specifically, our empirical analysis suggests that semantically related images tend to originate from nearby points in the noise space.
Building on these observations, we propose leveraging a dynamical systems approach that has not yet been explored in the context of distillation: Koopman operator theory. This theory is a classical yet recently revitalized approach for representing nonlinear systems with linear dynamics in an embedded space koopman1931hamiltonian ; rowley2009spectral . The core idea is that by mapping the system into a suitable embedding space, its evolution becomes linear. Thus, learning the diffusion process then reduces to jointly learning these embeddings and the linear operator. While Koopman operators are typically infinite-dimensional, we prove that an exact finite-dimensional representation exists under mild assumptions iacob2023finite . Moreover, prior work suggests that the Koopman operator preserves the structure inherent to certain dynamical systems goyal2025deep ; boulle2024multiplicative . This motivates the use of Koopman theory to harness the observed coherent semantic structure in diffusion dynamics, offering a principled foundation for accurate teacher model distillation. To this end, we formally connect the observed structure in diffusion models to the Koopman framework by proving that this structure is preserved within the learned dynamical representation.
[…]
Figure 1: A Gaussian sample (bottom left) evolves into a clean image (bottom right) via nonlinear dynamics $\Phi_{T}$. Leveraging the Koopman framework, learned encoders $E_{\varphi}$ and $E_{\phi}$ transform noise and image into an embedding space where the evolution becomes linear under $C_{\mu}$ (top).
In practice, we propose KDM (Koopman Distillation Model), a framework for modeling diffusion processes through linear evolution in a learned embedding space. Given a noisy input, an encoder maps it into an embedding space where a learned Koopman operator evolves it forward, and a decoder reconstructs the clean sample. A schematic illustration of our framework is presented in Fig. 1. The training objective encourages semantic preservation (via reconstruction loss), accurate latent linear dynamics (via Koopman consistency), and fidelity in output space (via prediction and adversarial losses). This design naturally yields fast, single-step generation, while preserving structural coherence distillation. We evaluate our method on the offline distillation benchmark, while adding two higher-resolution datasets, and achieve highly competitive results in both unconditional and conditional generation. These results demonstrate the effectiveness of our Koopman-based formulation in bridging the performance gap with online methods, offering a fast and high-fidelity offline diffusion model distillation. Our key contributions are:
  1. 1.
    Observing diffusion dynamics. We identify, analyze and formalize two key observations: (i) tools from dynamical systems theory, specifically Koopman operator theory, remain underexplored in diffusion modeling and distillation; (ii) training induces coherent semantic organization in noise space. These insights motivate a structure-promoting dynamics-aware approach to distillation.
  2. 2.
    A principled and practical Koopman-based distillation framework. We develop a theoretical foundation proving the existence of finite-dimensional Koopman representations and semantic structure preservation under mild assumptions and introduce a simple, scalable encoder–linear-dynamics–decoder architecture that preserves the teacher generation dynamics semantic structure.
  3. 3.
    Highly competitive results. We conduct a comprehensive unconditional and conditional evaluation showing that our method outperforms prior offline distillation approaches, achieving FID improvement, while enabling efficient, single-step generation.
2 Related Work
Generative Modeling.
Deep generative models have made significant progress in recent years, with diffusion models emerging as a leading framework for high-quality image, audio, and molecular generation. Building on principles from nonequilibrium thermodynamics and score matching sohl2015deep ; song2021score , these models define a forward process that gradually corrupts data with noise and learn to reverse it via a neural network-based denoising process. Variants like DDPMs ho2020denoising and score-based generative models song2021score have shown competitive performance across a wide range of benchmarks, often surpassing VAEs and GANs kingma2014auto ; goodfellow2014generative in sample fidelity and diversity.
Accelerating Diffusion Sampling.
Diffusion models generate high-quality samples but are slow due to iterative sampling. One approach improves integration using higher-order solvers and expressive noise schedules song2020denoising ; lu2022dpm , achieving high-quality results in significantly fewer steps. Another class of approaches seeks to distill a student model from a pre-trained teacher, reducing the sampling process to a single or few steps. Online distillation techniques luhman2021knowledge ; salimans2022progressive supervise the student directly with teacher predictions during training, but require continuous access to the full teacher model, resulting in high memory and compute overhead. Recent methods demonstrate high-quality one-step generation across CIFAR-10, ImageNet, and text-to-image benchmarks yin2024one ; zhou2024score , using objectives based on distribution matching yin2024one , score identity zhou2024score , adversarial training sauer2024fast ; sauer2024adversarial , and consistency models song2023consistency ; kim2024consistency among others chen2023geometric ; luo2023diff ; nguyen2024swiftbrush ; zhou2024simple ; berman2025said . Offline distillation methods geng2023one rely on teacher-generated noise-data pairs and decouple student training from the teacher. Aside from the concurrent work turan2025unfolding , current distillation methods largely unexplored Koopman-based perspectives and their relation to the teacher’s underlying dynamics.
Koopman-Based Modeling.
Koopman operator theory provides a powerful framework for modeling nonlinear dynamical systems via linear operators acting on observable functions koopman1931hamiltonian ; rowley2009spectral . This view has been applied to tasks such as deep learning of dynamical systems takeishi2017learning ; lusch2018deep ; yeung2019learning ; otto2019linearly ; naiman2023operator ; naiman2024generative , sequential disentanglement berman2023multifactor ; barami2025disentanglement , and control han2022desko . A related strand of work focuses on identifying coherent sets—regions of the state space that evolve coherently over time—using Koopman-based embeddings mezic2005spectral ; froyland2010coherent . While these works exploit the Koopman structure for interpretability or stability, our work focuses on using Koopman-based modeling for distilling diffusion models. In this context, we demonstrate that the learned embedding space in diffusion models encodes semantically meaningful structure, and we leverage this to design a principled offline distillation framework.
3 Mathematical Background
Deterministic Diffusion Sampling.
We consider deterministic sampling in diffusion models, as introduced in the EDM framework karras2022elucidating . A pre-trained diffusion model defines a time-indexed denoising vector field $f_{\theta}:\mathbb{R}^{n}\times[0,T]\rightarrow\mathbb{R}^{n}$, which maps a noisy latent state $x_{t}\in\mathbb{R}^{n}$ at time $t\in[0,T]$ toward the data manifold (at time 0). Sampling is performed by solving the reverse-time ordinary differential equation (ODE),
$\frac{\operatorname*{d}x_{t}}{\operatorname*{d}t}=f_{\theta}(x_{t},t)\ ,\quad x_{T}\sim\mathcal{N}(0,\sigma_{T}^{2}\mathbf{I})\ ,$ (1)
integrated backward from a Gaussian sample $x_{T}$ at time $T$ to a clean sample $x_{0}$ at time $t=0$. The function $f_{\theta}$ is typically trained to approximate the score function $\nabla_{x_{t}}\log p_{t}(x_{t})$ or a reparameterization thereof, such as the probability flow song2021score . These deterministic formulations enable sampling via high-order ODE solvers and provide a tractable setting for distillation and trajectory analysis karras2022elucidating .
Koopman Operator Theory.
Let $\Phi_{t}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ define a (possibly time-dependent) dynamical system mapping states forward in time, i.e., $x_{t}=\Phi_{t}(x_{0})$. The Koopman operator $\mathcal{K}_{t}$ koopman1931hamiltonian is a linear operator that acts on scalar-valued observables $g:\mathbb{R}^{n}\rightarrow\mathbb{C}$ via composition,
$\mathcal{K}_{t}g:=g\circ\Phi_{t}\ ,$ (2)
mapping each observable to its pullback along the dynamics. While the original system $\Phi_{t}$ may be nonlinear, $\mathcal{K}_{t}$ is linear on the space of observables $L^{2}(\mathbb{C}^{n})$ (with respect to a suitable measure) eisner2015operator . In classical Koopman theory, this operator is infinite-dimensional. However, if there exists a finite-dimensional subspace $\mathcal{O}=\text{span}\{g_{1},\dots,g_{d}\}\subset L^{2}(\mathbb{C}^{n})$ that is invariant under $\mathcal{K}_{t}$ for all $t\in[0,T]$, then the dynamics of $\Phi_{t}$ can be exactly represented in this lifted space via a family of linear operators $C_{t}\in\mathbb{R}^{d\times d}$, $t\in[0,T]$, such that for all $g\in\mathcal{O},\,\mathcal{K}_{t}g=C_{t}g$. This allows modeling nonlinear dynamics through linear evolution in a learned observable space, a principle that underpins our Koopman-based distillation framework.
4 Koopman Distillation of Diffusion Models
We propose KDM, a principled offline distillation framework for diffusion models based on Koopman operator theory. The key idea is to model the dynamics of the generative process, specifically the transformation from a noisy sample $x_{T}$ to its clean counterpart $x_{0}$, as a single-step linear evolution in a learned embedding space. This approach enables fast and structure-preserving generation in a single forward pass, without requiring access to the full diffusion trajectory or teacher model at training time.
Dynamical System Perspective on Diffusion Models.
We build upon two key observations regarding diffusion models. First, diffusion can be understood as a dynamical system, where the evolution of a noisy data point $x_{t}$ is governed by a time-dependent velocity field, see Eq. (1). This continuous-time formulation admits an integrated flow map $\Phi_{t}$ such that $x_{t}=\Phi_{t}(x_{0})$, where $\Phi_{t}$ evolves the initial condition $x_{0}$ forward in time. This formulation makes it natural to propose adopting a Koopman operator perspective, a novel approach in the context of distillation, which enables the lifting of nonlinear dynamics into a linear evolution in function space under suitable choices of observables, see Eq. (2).
The second observation considers the structure of the trained diffusion dynamics. Score‑based diffusion models learn the score $\nabla_{x_{t}}\log p_{t}(x_{t})$ of noise‑corrupted data at multiple noise levels and then reverse that specific corruption process using Langevin dynamics song2021score or learned reverse kernels ho2020denoising to transform Gaussian noise into images. This framework has been unified through the lens of stochastic differential equations (SDEs) by song2021score . Flow matching models lipman2023flow learn a continuous velocity field that transports a simple reference distribution directly into the data distribution, where sampling is typically performed by integrating the learned ordinary differential equation (ODE). In what follows, we study the structure of the dynamical mapping arising from diffusion-based learning.
We pose the following question: Is there semantic structure in the generative process of diffusion models? In general, prior to training, there is no assurance of semantic structure between the noise and data distributions (see Fig. 2, left). However, after training, we observe the emergence of a coherent and semantically meaningful mapping from the Gaussian noise distribution to the data distribution (see Fig. 2, right). Specifically, by utilizing a simple 2D checkerboard toy dataset, we train an EDM karras2022elucidating model and generate 50,000 samples, recording both the initial noise vectors and their corresponding generated outputs on the checkerboard. We then color-code each sample by its target checker cell and assign the same color to its originating noise vector. This visualization reveals that samples originating from the same semantic region in data space tend to cluster together in noise space, indicating structured and coherent dynamics. Notably, we observe a similar pattern in both flow matching model and our KDM. This emergence of coherent structure, together with prior evidence that the Koopman operator preserves structure in certain dynamical systems goyal2025deep ; boulle2024multiplicative , highlights the relevance of Koopman theory to diffusion dynamics. Indeed, our theoretical analysis (Sec. 5) establishes a formal link between this observed structure and Koopman theory, motivating its adoption as a principled framework for modeling these dynamics.
[…]
Figure 2: Untrained noise distributions are semantically unstructured (left). Training EDM on the data reveals the emergence of coherent clusters in noise space, i.e., at time $T$ (right).
Koopman-based Distillation.
Our Koopman distillation framework is designed by the following principle: given pairs of noisy and clean data $(x_{T},x_{0})$, where $x_{0}=\Phi_{T}^{-1}(x_{T})$ for some unknown nonlinear map $\Phi_{t}^{-1}$, we seek to model this transformation using a finite-dimensional linear operator in an embedding space, following Koopman theory. To achieve this, we introduce observable functions $\{g_{1},g_{2},...,g_{d}\}$, parameterized as encoder networks $E_{\phi},E_{\varphi}$ for $x_{0},x_{T}$, respectively, that map states from the original data space into a latent observable space. In this space, we assume the dynamics can be modeled linearly by a finite-dimensional Koopman operator linear layer, $C_{\eta}$. Finally, a corresponding decoder network $D_{\psi}$ maps the evolved observables back to the data space. The overall modeling objective becomes:
$x_{0}\approx D_{\psi}(C_{\eta}\,E_{\varphi}(x_{T}))\ .$ (3)
This formulation distills the potentially complex, nonlinear generative process into a simple one-step linear map in the embedded observables space, while maintaining fidelity to the original dynamics.
Koopman’s Loss Functions.
In what follows, $d(\cdot,\cdot)$ denotes a suitable distance function, such as mean squared error (MSE) or a perceptual loss (e.g., LPIPS zhang2018unreasonable ). To effectively learn the components $E_{\phi}$, $D_{\psi}$, and $C_{\eta}$, we define a composite loss comprising three terms (see similar objectives in lusch2018deep ; azencot2020forecasting ):
$\mathcal{L}_{\text{Koopman}}=\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{lat}}+\mathcal{L}_{\text{pred}}\ .$ (4)
The first term, the reconstruction loss $\mathcal{L}_{\text{rec}}$, ensures that the encoder-decoder pair $(E_{\phi},D_{\psi})$ can faithfully reconstruct $x_{0}$:
$\mathcal{L}_{\text{rec}}=d\big(x_{0},D_{\psi}(E_{\phi}(x_{0})\big)\ .$ (5)
This term regularizes the embedding by encouraging $E_{\phi}$ to preserve sufficient semantic information about the data, while not encoding any dynamic features. The second term, the latent dynamics loss $\mathcal{L}_{\text{lat}}$, enforces consistency of the linear evolution in the observable space:
$\mathcal{L}_{\text{lat}}=d\big(E_{\phi}(x_{0}),C_{\eta}E_{\varphi}(x_{T})\big)\ ,$ (6)
where $C_{\eta}$ is the Koopman operator in latent space. This term ensures that applying $C_{\eta}$ to the embedded noisy state $E_{\varphi}(x_{T})$ approximates the embedded clean state $E_{\phi}(x_{0})$. $C_{\eta}$ is either implemented with a regular learning linear layer or by a new approach we develop that factorizes the matrix, which enables eigenspectrum control (KDM-F). We describe the decomposition and its merits in App. C.
Finally, the prediction loss $\mathcal{L}_{\text{pred}}$ evaluates the full end-to-end prediction after encoding, applying the Koopman operator, and decoding:
$\mathcal{L}_{\text{pred}}=d\big(x_{0},D_{\psi}(C_{\eta}E_{\varphi}(x_{T}))\big)\ ,$ (7)
encouraging the system to accurately predict $x_{0}$ from $x_{T}$ in a single forward pass through the Koopman pipeline. Together, these terms provide complementary supervision signals.
Adversarial Loss.
Adversarial frameworks are widely employed in distillation methods sauer2023adversarial ; kim2024consistency ; ladd2024latent ; sdxllightning2024progressive . For example, kim2024consistency leverages an adversarial loss to enhance the fidelity of one-step generation. Motivated by these approaches, we introduce an auxiliary adversarial loss to further refine the distillation process. Specifically, we introduce a discriminator network $D_{\gamma}$, parameterized by $\gamma$, and define the adversarial loss as:
$\mathcal{L}_{\text{adv}}(\phi,\psi,\gamma)=\mathbb{E}_{x_{0}\sim p_{\text{data}}}[\log D_{\gamma}(x_{0})]+\mathbb{E}_{x_{T}\sim p_{\text{data}}}[\log(1-D_{\gamma}(\hat{x}_{0}))]\ ,$ (8)
where $\hat{x}_{0}=D_{\psi}(C_{\eta}E_{\varphi}(x_{T}))$ is the Koopman-predicted sample. Thus, the final training objective is:
$\mathcal{L}=\mathcal{L}_{\text{Koopman}}+\lambda_{\text{adv}}\mathcal{L}_{\text{adv}}\ ,$ (9)
where $\lambda_{\text{adv}}$ balances the adversarial component. In practice, throughout our experiments, we choose $\lambda_{\text{adv}}=0.01$. Importantly, we utilized a simple $D_{\gamma}$ with four convolutional layers. By enhancing the perceptual quality of the predictions without altering the fundamental Koopman modeling, this adversarial augmentation strengthens the overall generation performance while preserving the structured distillation at the heart of our method blau2018perception . We present losses and more ablation studies in App. D.1. Additionally, A pseudo-code of our KDM is provided in App. E.1.
Control Modeling for Conditional Dynamics.
A crucial aspect of generative modeling with denoising models is the ability to condition generation on auxiliary signals or labels dhariwal2021diffusion . Koopman theory provides a natural framework to incorporate such conditioning through control signals that influence the Koopman operator azencot2013operator ; brunton2022modern . Formally, let $c$ denote the control signal or conditioning variable (e.g., a class label) that we want the generative process to be conditioned on. Our goal is to model the conditional dynamics $\Phi(x_{T},c)$. In the Koopman framework, we extend the observable and evolution operators to incorporate $c$, and formulate the generation as:
$D_{\psi}\big(C_{\eta}E_{\phi}(x_{T},c)+C_{\mu}(c),\,c\big)\approx\Phi(x_{T},c),$
where $C_{\mu}(c)$ is a control-dependent modulation term (e.g., a linear transformation of $c$). This formulation enables our model to adapt its generative trajectory based on the conditioning signal, seamlessly integrating the conditioning mechanism into the Koopman dynamics. We apply this control modeling in our conditional generation experiments.
5 Theoretical Properties of KDM
We now state the first key theoretical result, establishing the existence of an exact finite-dimensional Koopman representation under mild conditions. Specifically, for an analytic dynamical system, it is possible to approximate its nonlinear evolution to arbitrary accuracy by a finite-dimensional linear operator acting on lifted coordinates. In particular, a small number of polynomial observables can capture the system’s dynamics with controllable approximation error. This result is significant because it means we can faithfully replicate the complex behavior of the teacher model using a simpler, linear system in a transformed space. In practical terms, it ensures our student can learn to follow the teacher’s dynamics with high accuracy.
Theorem 5.1 (Finite Koopman Operator for Analytic Dynamics).
Let $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be an analytic map, and let $x_{T}\sim\mathcal{N}(0,I_{n})$. Then, for any $\epsilon>0$, there exists a linear operator $C\in\mathbb{R}^{d\times d}$ such that
$\mathbb{E}\left[\|\xi(\Phi(x_{T}))-C\xi(x_{T})\|^{2}\right]\leq\epsilon\ ,$ (10)
where $\xi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ denotes the observable lifting map. Moreover, the required dimension $d$ grows at most polynomially with $1/\epsilon$, and the mapping $\Phi$ can be approximated arbitrarily well using multivariate polynomials up to degree $d$.
The proof builds on the analytic formulation of $\Phi$, allowing a convergent multivariate Taylor expansion trefethen2019approximation , and on exact finite-dimensional Koopman operators for polynomial dynamics iacob2023finite , see App. A.
Our second theoretical result concerns the structure induced by the reverse diffusion dynamics, under a set of assumptions established by Thm. 5.1. Under these conditions, the theorem states that proximity in the Koopman coordinate representation at the initial noise time implies semantic similarity between the corresponding samples at data time. In practice, the theorem says that if two inputs start out close together in the transformed space, the images they generate will be semantically similar. This provides a theoretical foundation for the structure-preserving behavior of our model, as observed empirically in Sec. 6.1.
Theorem 5.2 (Semantic Proximity via Koopman-Invariant Coordinates).
Let $\Phi_{t}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the reverse diffusion flow of a trained model (from time $T$ to 0), and let $\mathcal{O}\subset L^{2}(\mathbb{R}^{n})$ be a finite-dimensional subspace such that:
  1. 1.
    $\mathcal{K}_{t}\mathcal{O}\subset\mathcal{O}$ for all $t\in[0,T]$,
  2. 2.
    $\exists\,C_{t}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\mathcal{K}_{t}\xi=\xi\circ\Phi_{t}=C_{t}\xi$ for all $\xi\in\mathcal{O}$,
  3. 3.
    $\Phi_{T}$ is locally Lipschitz.
For any $x_{T}^{1},x_{T}^{2}\in\mathbb{R}^{n}$, define $x_{0}^{j}:=\Phi_{T}(x_{T}^{j})$ for $j=1,2$, and let $\xi:\mathbb{R}^{n}\to\mathbb{R}^{d}$ be the vector-valued observable $\xi(x):=(\xi_{1}(x),\dots,\xi_{d}(x))\in\mathcal{O}$. Then:
$\|x_{0}^{1}-x_{0}^{2}\|\leq L\cdot\|\Phi_{T}(x_{T}^{1})-\Phi_{T}(x_{T}^{2})\|\leq L\cdot\|C_{T}\|\cdot\|\xi(x_{T}^{1})-\xi(x_{T}^{2})\|.$
The proof builds on the observation that diffusion model dynamics tend to preserve semantic similarity between inputs and outputs. By leveraging the Koopman operator’s spectral properties, we construct a finite-dimensional space of functions that remains stable under the model’s evolution. Within this space, the Koopman coordinates $\xi$ provide a locally bi-Lipschitz representation, while the reverse diffusion map $\Phi_{T}$ is Lipschitz continuous. Together, these properties ensure that semantically similar inputs remain close throughout the diffusion process. Note that the theorem holds also for $\xi:=\text{id}$, i.e., the identity map mauroy2020introduction ; dogra2020optimizing . The full proof is given in App. B.
Importantly, we observe in practice an average spectral norm of the Koopman operator $C_{T}$ of $0.88\pm 0.18$, indicating a reasonably tight bound and supporting the practical utility of our theory. Moreover, our framework enables explicit control over this norm via eigenvalue penalties on the Koopman operator. As detailed in App. C, we factorize $C_{T}$ to permit direct constraints on its spectrum (e.g., radius or banding), providing a principled regularizer on $\|C_{T}\|$.
6 Experiments
We begin by further investigating the emergent dynamics structure that map noise to images, and its presence in KDM (Sec. 6.1). We then extensively assess our method on the offline distillation benchmark (Sec. 6.2), including analyses of computational complexity, data and parameter scaling, and denoiser-agnostic utility. Additionally, an ablation study of the loss functions and adversarial losses stability is provided in App. D.1. Our code is in https://github.com/azencot-group/KDM.
Setup and evaluation.
We follow the protocol of GET geng2023one for generating noisy-clean pairs and for evaluation, using the CIFAR-10 krizhevsky2009learning , FFHQ $\times64$ karras2019style , and AFHQv2 $\times64$ choi2020stargan datasets. Our training setup matches that of PD salimans2022progressive and GET geng2023one . For evaluation, we report image quality using Fréchet Inception Distance (FID) and Inception Score (IS), computed over 50k samples. Both encoders $E_{\varphi}$ and $E_{\phi}$, as well as the decoder $D_{\psi}$, are implemented as compact UNets with reduced output channels to ensure a similar parameter size to the other single-step methods. The discriminator consists of four simple Conv2D layers. Full implementation details are provided in App. E.3.
6.1 Observations of Emergent Latent Structure
$\sigma{=}$0 0.05 0.1 0.2 0.3 0.4 0.5 0.6 $\sigma{=}$0 0.05 0.1 0.2 0.3 0.4 0.5 0.6
Refer to caption Refer to caption
Figure 3: Visualization of vicinity comparison across different noise levels $\sigma$. Left: EDM karras2022elucidating . Right: our method. Each image corresponds to a small perturbation around the original noise sample.
Real-world data dynamics: local structure analysis.
Extending the 2D toy experiment from Sec. 4, we now explore the high-dimensional FFHQ $\times64$ dataset. Since dimensionality reduction can obscure fine-grained patterns, we instead analyze the local neighborhood of a single latent sample $x_{T}\sim\mathcal{N}(0,I_{n})$. We generate perturbed versions $\hat{x}_{T}=x_{T}+\varepsilon\cdot\sigma$ with $\varepsilon\sim\mathcal{N}(0,I_{n})$, and normalize each $\hat{x}_{T}$ using its own mean and standard deviation. This setup tests whether semantic structure emerges locally around $x_{T}$. We vary $\sigma\in\{0,0.05,\dots,0.60\}$ and visualize the outputs in Fig. 3 (left). For all $\sigma<0.3$, the images remain semantically aligned with the original image (when $\sigma=0$), indicating local smoothness and manifold adherence. As $\sigma$ grows, semantics diverge but retain global coherence even at $\sigma=0.60$. The same experiment on KDM (Fig. 3, right) reveals similar behavior, affirming that KDM preserves the local semantic structure of the original EDM. Our results show that both models organize latent space into meaningful neighborhoods, even in high-dimensional settings.
Visual analysis of dynamics distillation.
Beyond preserving local structure, KDM also approximately reproduces the noise-to-image mapping of EDM. When starting from the same noise vector $x_{T}\sim\mathcal{N}(0,I_{n})$, sampled using seeds not seen during training, both EDM (with 79 NFEs) and KDM (with a single NFE) generate perceptually similar outputs (see Fig. 4). This suggests that KDM generalizes well beyond its training data and effectively distills the mapping from noise to data. We observe this consistency across multiple datasets and include further qualitative results in App. D.3. To quantify this, we generate 50k samples using both methods and compare outputs using LPIPS and MSE. KDM achieves scores of 0.13 (LPIPS) and 0.009 (MSE), in contrast to 0.48 and 0.57 when comparing randomly permuted images, demonstrating the similarity between the models’ generation. To complement quantitative metrics, we conduct a human study where participants choose the more realistic image collage between our model and EDM (Fig. 4; details in App. E.4). Evaluators showed no clear preference: $55\%$ of 300 choices favored our model, highlighting its strong fidelity.
Figure 4: Sampled images using 79 NFEs with EDM karras2022elucidating (left), and using 1 NFE with KDM (right).
Outliers and their relation to the dynamics structure.
We analyze a recurring failure mode in both the teacher and distilled models: the generation of outlier samples that fall outside the designated checkerboard regions. These out-of-distribution points often appear in the final outputs and reflect a limitation in capturing the intended data distribution. To investigate their origin, we trace these outliers back through the generation process and identify two primary sources: (1) the tails of the Gaussian prior and (2) regions near decision boundaries between adjacent checkerboard cells. Fig. 5C shows this behavior for the distilled model (black crosses denote outliers); additional comparisons with the teacher are provided in App. E.7. These findings suggest that the model’s dynamics are more error-prone in high-uncertainty regions, offering potential directions for improvement. Integrating more robust sampling techniques or leveraging self-supervised signals may help mitigate such failures. This insight could also inform future efforts in uncertainty estimation.
[…]
Figure 5: A) Model size vs FID. B) Data Size vs FID. C) Outlier analysis
6.2 One-step Generation Results
Unconditional and conditional generation with CIFAR-10.
We evaluate on the standard offline distillation benchmark of (geng2023one, ), which distills a pre-trained EDM on CIFAR-10. Whereas the EDM teacher samples with 35 function evaluations (NFEs), our student attains comparable quality with a single NFE. We also adapt two recent SOTA baselines, IMM (zhou2025inductive, ) and RF-2++ (lee2024improving, ), to the same strictly offline setting, where only teacher-generated data are available and the original model weights (and online queries) are not accessible (the RF-2++ variant corresponds to its “w/o pre-train” configuration). All methods are evaluated under identical protocols. Tab. 3 reports unconditional generation results, where our method achieves highly competitive performance, surpassing the previous offline baseline in both FID and IS. In particular, we observe $\approx 25\%$ reduction in FID over GET. Tab. LABEL:table:conditional-cifar shows results for class-conditional generation, where our model yields $\approx 40\%$ improvement in FID and superior IS compared to GET. Although offline distillation offers advantages in efficiency, privacy, and flexible deployment fernandez2023privacy ; geng2023one , a performance gap with online methods remains. This may partly be because online approaches leverage the full generation trajectory, whereas, in the offline pairwise setup, methods only observe two endpoints or just the generated data, placing them at a natural disadvantage. Our results demonstrate a substantial step toward closing this gap.
Furthermore, unlike prior methods that rely on architectural assumptions geng2024consistency ; lee2024improving , our approach is denoiser-agnostic and successfully distills both EDM and flow matching (FM) models (see Tab. 3), demonstrating broad applicability. Finally, we implemented a factorized KDM (KDM-F). Since every matrix $C_{\eta}$ is diagonalizable up to an arbitrarily small perturbation of the entries axler2015linear , we can write: $C_{\eta}=P\Lambda P^{-1}$ where $P\in\mathbb{C}^{d\times d}$ is an invertible matrix and $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})\in\mathbb{C}^{d\times d}$. While enabling efficient implementation of spectral penalties on $\Lambda$ (see complexity analysis below), this approach preserves strong performance and accuracy.
Table 1: Generative performance on unconditional CIFAR-10.
Models $(\downarrow$) / Metrics ($\rightarrow$) NFE $\downarrow$ FID $\downarrow$ IS $\uparrow$
Diffusion Models
DDPM ho2020denoising 1000 3.17 9.46
Score SDE song2020denoising 2000 2.2 9.89
EDM karras2019style 35 2.04 9.84
Continuous Flows
Flow Matching (Diffusion) lipman2022flow 183 8.06 -
Flow Matching (OT) tong2023improving 50 4.94 -
2-rf++ lee2024improving 1 3.07 -
GANs
StyleGAN-XL karras2020training 1 1.85 -
Diffusion Distillation
PD salimans2022progressive 1 9.12 -
DFNO zheng2023fast 1 4.12 -
iCD-EDM (LPIPS) song2023improved 1 2.83 9.54
CTM (LPIPS) song2023improved 1 1.98 -
ECM geng2024consistency 1 4.54 -
Offline Distillation
GET geng2023one 1 6.91 9.16
2-rf++ w/o pre-train lee2024improving 1 6.32 9.01
IMM (zhou2025inductive, ) 1 4.81 9.16
KDM (EDM) 1 4.65 9.21
KDM (FM) 1 5.91 8.86
KDM-F (EDM) 1 4.68 9.08
Table 2: Generative performance on class-conditional CIFAR-10.
Models $(\downarrow$) / Metrics ($\rightarrow$) NFE $\downarrow$ FID $\downarrow$ IS $\uparrow$
GANs
BigGAN brock2018large 1 14.73 9.22
StyleGAN2-ADA karras2020training 1 2.42 10.14
Diffusion Distillation
Guided Dist. (w=0.3w=0.3meng2023distillation 1 7.34 8.90
ECM geng2024consistency 1 3.81 -
Offline Distillation
GET geng2023one 1 6.25 9.40
KDM (EDM) 1 3.56 9.54
KDM-F (EDM) 1 3.24 9.68
Table 3: Generative performance on unconditional generation on FFHQ $\times64$ and AFHQv2 $\times64$. *Indicates our reproduction.
Models $(\downarrow$) / Metrics ($\rightarrow$) NFE $\downarrow$ FID $\downarrow$ IS $\uparrow$
FFHQ $\times64$
EDM* 79 2.47 3.37
KDM (EDM) 1 6.54 3.12
AFHQv2 $\times64$
EDM* 79 2.02 9.04
KDM (EDM) 1 4.85 8.25
Unconditional generation with FFHQ and AFHQv2.
[…]
Scalability in parameters and data.
Offline distillation inherently supports architectural independence from the teacher model, allowing flexible scaling of the student to meet diverse deployment needs geng2023one ; geng2024consistency . To assess this, we evaluate our method against GET across three model sizes—20M, 38M, and 84M parameters (Fig. 5A). Our method consistently outperforms GET in FID scores at all scales, with especially notable gains at higher capacities (e.g., 4.89 vs. 7.19 at 84M), demonstrating its effectiveness for both lightweight and high-capacity settings. In parallel, we investigate how performance scales with training data size, a key factor in offline distillation. Following GET geng2023one , we train on 250K, 500K, and 1M samples and report FID scores on unconditional CIFAR-10 in Fig. 5B. Results show that more data yields better generation quality. For instance, increasing the dataset from 250K to 1M samples results in a $\approx 33\%$ FID reduction, showing the importance of data scale. While our method remains data-efficient, additional data can enhance performance.
Complexity analysis.
Offline distillation can offer advantages over online methods, both memory-wise and time-wise (see App. E.5). KDM further improves over GET. As shown in Tab. 4, it matches GET in parameter count but achieves a $4\times$ speedup per training iteration and over $8\times$ faster sampling. Notably, KDM-F (App. C) maintains this efficiency even with spectral regularization. Specifically, we evaluate KDM and KDM-F with an additional eigen-loss (EL) term, observing that KDM-F + EL requires training time comparable to KDM (without EL), whereas KDM + EL incurs a longer training duration. These results underscore the practical scalability of our approach.
Table 4: Parameters in millions, time in seconds. RTX 4090 GPU used with batch size 128.
Metric GET KDM KDM + EL KDM-F + EL
Params (M) 62 62 62 65
Train (s) 3.60 0.91 4.82 0.97
Sampling (s) 937 114 114 116
7 Discussion
In this work, we present two key observations. First, we reinterpret the distillation process through the lens of dynamical systems theory. Second, we uncover the semantical structure in the mapping from latent space to data space in trained diffusion dynamics. Motivated by these insights, we propose a Koopman-based framework for distilling diffusion models, bridging operator theory and generative modeling. We theoretically establish the existence of a Koopman operator that captures the underlying generative process, and demonstrate, both theoretically and empirically, that our method preserves its structure. This perspective also offers insights into potential failure modes. Beyond its conceptual contributions, our approach is practical and efficient, achieving highly competitive single-step offline distillation. We further analyze its robustness and scalability. We hope this work lays the foundation for future directions, including leveraging richer training signals and extending distillation techniques to advanced text-to-image generative models. See App. F for further discussion.
Acknowledgments
This research was partially supported by the Lynn and William Frankel Center of the Computer Science Department, Ben-Gurion University of the Negev, an ISF grant 668/21, an ISF equipment grant, and by the Israeli Council for Higher Education (CHE) via the Data Science Research Center, Ben-Gurion University of the Negev, Israel.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Contributions are clearly state at the end of the intro
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We dedicate an discussion in the appendix for improvements and limitations
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide in the appendix full proofs
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide technical details in the appendix and code in the supplementary
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Yes, in the supplementary material we provide code and link to anonymous drive with the data
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The code provided reconstructs the exact experimental setting
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All results are reported by the standard protocol for generative models evaluation
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide complexity analysis and hardware information.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [N/A]
  44. Justification:
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification:
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification:
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [N/A]
  59. Justification:
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: See codebase
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification:
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification:
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification:
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Proof of Theorem 5.1
For the reader’s convenience, we restate the theorem.
Theorem (Finite Koopman Approximation for Analytic Dynamics).
Let $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be an analytic map, and let $x_{T}\sim\mathcal{N}(0,I_{n})$. Then, for any $\epsilon>0$, there exists a linear operator $C\in\mathbb{R}^{d\times d}$ such that
$\mathbb{E}\left[\|\xi(\Phi(x_{T}))-C\xi(x_{T})\|^{2}\right]\leq\epsilon\ ,$
where $\xi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ denotes the observable lifting map. Moreover, the required dimension $d$ grows at most polynomially with $1/\epsilon$, and the mapping $\Phi$ can be approximated arbitrarily well using multivariate polynomials up to degree $d$.
Proof.
Since $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is analytic, for every $\delta>0$ and every compact set $K\subset\mathbb{R}^{n}$, there exists a multivariate polynomial map $p:\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that
$\sup_{x\in K}\|\Phi(x)-p(x)\|\leq\delta\ ,$
which follows from the theory of approximation by Taylor polynomials for analytic functions trefethen2019approximation .
Let $x_{T}\sim\mathcal{N}(0,I_{n})$. Since $x_{T}$ has exponentially decaying tails vershynin2018high , for any $\eta>0$ there exists $R>0$ such that
$\mathbb{P}(x_{T}\notin B_{R})\leq\eta,$
where $B_{R}=\{x\in\mathbb{R}^{n}:\|x\|\leq R\}$. Restricting attention to $B_{R}$, we consider the polynomial approximation $p$ of $\Phi$ with error at most $\delta$.
Now, following the result of Iacob et al. iacob2023finite , for the polynomial map $p$, there exists a finite-dimensional lifting map $\xi:\mathbb{R}^{n}\to\mathbb{R}^{d}$ consisting of monomials up to some degree, and a linear operator $C\in\mathbb{R}^{d\times d}$, such that
$\xi(p(x))=C\xi(x)$
for all $x\in\mathbb{R}^{n}$. In particular, on $B_{R}$,
$\|\xi(\Phi(x))-C\xi(x)\|\leq\|\xi(\Phi(x))-\xi(p(x))\|,$
since $\xi(p(x))=C\xi(x)$ exactly.
Since $\xi$ is a smooth (polynomial) map, there exists a Lipschitz constant $L>0$ depending on the derivatives of $\xi$ over $B_{R}$, such that
$\|\xi(\Phi(x))-\xi(p(x))\|\leq L\|\Phi(x)-p(x)\|\leq L\delta$
for all $x\in B_{R}$.
Thus, for $x_{T}\in B_{R}$,
$\|\xi(\Phi(x_{T}))-C\xi(x_{T})\|\leq L\delta.$
Outside $B_{R}$, the trivial bound
$\|\xi(\Phi(x_{T}))-C\xi(x_{T})\|\leq 2M$
holds for some finite constant $M>0$, due to the polynomial growth of $\xi$ and the Gaussian decay of $x_{T}$.
Hence, the expected Koopman error satisfies
$\mathbb{E}\left[\|\xi(\Phi(x_{T}))-C\xi(x_{T})\|^{2}\right]\leq(1-\eta)L^{2}\delta^{2}+\eta(2M)^{2}.$
Given any $\epsilon>0$, we can choose $\eta$ sufficiently small so that $\eta(2M)^{2}\leq\epsilon/2$, and then $\delta$ sufficiently small so that $(1-\eta)L^{2}\delta^{2}\leq\epsilon/2$.
Notably, the degree of the polynomial map $p$ needed, and hence the number of monomials $d$ in $\xi$, grows at most polynomially with $1/\delta$ and thus with $1/\epsilon$, concluding the proof.
We further note that the function $\Phi$ implemented by DhariwalUNet karras2022elucidating is analytic. Specifically, each module composing the network—including convolutions, SiLU activations, residual connections, attention mechanisms, and group normalization with $\epsilon>0$—is an analytic function. Since compositions of analytic functions remain analytic, it follows that the entire forward pass $\Phi$ is analytic on $\mathbb{R}^{n}$.
Appendix B Proof of Theorem 5.2
We restate the theorem for the readers’ convenience.
Theorem (Semantic Proximity via Koopman-Invariant Coordinates).
Let $\Phi_{t}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the reverse diffusion flow of a trained model (from time $T$ to 0), and let $\mathcal{O}\subset L^{2}(\mathbb{R}^{n})$ be a finite-dimensional subspace such that:
  1. 1.
    $\mathcal{K}_{t}\mathcal{O}\subset\mathcal{O}$ for all $t\in[0,T]$,
  2. 2.
    $\exists\,C_{t}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\mathcal{K}_{t}\xi=\xi\circ\Phi_{t}=C_{t}\xi$ for all $\xi\in\mathcal{O}$,
  3. 3.
    $\Phi_{T}$ is locally Lipschitz.
For any $x_{T}^{1},x_{T}^{2}\in\mathbb{R}^{n}$, define $x_{0}^{j}:=\Phi_{T}(x_{T}^{j})$ for $j=1,2$, and let $\xi:\mathbb{R}^{n}\to\mathbb{R}^{d}$ be the vector-valued observable $\xi(x):=(\xi_{1}(x),\dots,\xi_{d}(x))\in\mathcal{O}$. Then:
$\|x_{0}^{1}-x_{0}^{2}\|\leq L\cdot\|\Phi_{T}(x_{T}^{1})-\Phi_{T}(x_{T}^{2})\|\leq L\cdot\|C_{T}\|\cdot\|\xi(x_{T}^{1})-\xi(x_{T}^{2})\|.$
Proof.
Let $\Phi_{t}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ denote the reverse diffusion dynamics. Based on the proof of Thm. 5.1, there exists a linear operator $C_{t}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ satisfying
$\mathcal{K}_{t}\xi=\xi\circ\Phi_{t}=C_{t}\xi\ ,$
for every $\xi\in\mathcal{O}$ and $t\in[0,T]$.
Define the vector-valued map for the basis $\{\xi_{j}\}$ of $\mathcal{O}$
$\xi(x):=(\xi_{1}(x),\dots,\xi_{d}(x))\in\mathbb{R}^{d}.$
Since the $\xi_{j}$ are smooth and capture independent semantic directions, $\xi$ is locally bi-Lipschitz near typical $x_{T}$. That is, there exist constants $C_{1},C_{2}>0$ such that
$C_{1}\|\xi(x_{T}^{1})-\xi(x_{T}^{2})\|\leq\|x_{T}^{1}-x_{T}^{2}\|\leq C_{2}\|\xi(x_{T}^{1})-\xi(x_{T}^{2})\|.$
Because $\Phi_{T}$ is locally Lipschitz, there exists $L>0$ such that
$\|\Phi_{T}(x_{T}^{1})-\Phi_{T}(x_{T}^{2})\|\leq L\|x_{T}^{1}-x_{T}^{2}\|.$
Substituting the bi-Lipschitz inequality yields
$\|\Phi_{T}(x_{T}^{1})-\Phi_{T}(x_{T}^{2})\|\leq LC_{2}\|\xi(x_{T}^{1})-\xi(x_{T}^{2})\|.$
Moreover, by Koopman linearity,
$\xi(\Phi_{T}(x_{T}^{j}))=C_{T}\xi(x_{T}^{j}),\quad j=1,2,$
and therefore
$\xi(\Phi_{T}(x_{T}^{1}))-\xi(\Phi_{T}(x_{T}^{2}))=C_{T}(\xi(x_{T}^{1})-\xi(x_{T}^{2})).$
Thus, proximity in $\xi(x_{T})$ coordinates implies proximity of final samples $x_{0}^{1}$ and $x_{0}^{2}$ under $\Phi_{T}$.
Appendix C Factorized Koopman Matrix for Fast Eigenspecturm penalty
To enable control over eigenvalues of the Koopman operator, we describe an efficient implementation of our Koopman-based Distillation Method (KDM), which we term KDM-F, alleviating the need for computation of the eigendecomposition during training. Since every matrix $C:=C_{\eta}$ is diagonalizable up to an arbitrarily small perturbation of the entries axler2015linear , we can write:
$C=P\Lambda P^{-1}\,$
where $P\in\mathbb{C}^{d\times d}$ is an invertible matrix and $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})\in\mathbb{C}^{d\times d}$. This way enables efficient implementation of spectral penalties on $\Lambda$, without the need to compute it on the fly during training.
To represent this decomposition within a real-valued neural network framework, we learn the real and imaginary components of the Koopman eigenvectors separately. Specifically, two orthonormal matrices, $P_{\text{re}}$ and $P_{\text{im}}$, represent the real and imaginary parts of the eigenvectors. These are combined into a real-valued block matrix using the standard transformation for complex-to-real matrix representation:
$P=\begin{bmatrix}P_{\text{re}}&-P_{\text{im}}\\ P_{\text{im}}&P_{\text{re}}\end{bmatrix}\in\mathbb{R}^{2d\times 2d}\ .$
Rather than explicitly computing the inverse $P^{-1}$, we parameterize a separate learnable matrix $P_{\text{inv}}$ with the same structure, constructed analogously from two unconstrained real matrices. This improves flexibility and avoids numerical instability associated with matrix inversion.
The eigenvalues of the Koopman operator are encoded using polar coordinates. Specifically, the modulus and phase of each eigenvalue are parameterized as
$\lambda_{j}=e^{-\exp(\nu_{j})}\cdot e^{i\theta_{j}}=e^{-\exp(\nu_{j})}\left(\cos(\theta_{j})+i\sin(\theta_{j})\right),$
where $\nu_{j}$ and $\theta_{j}$ are learnable parameters. Let $\lambda_{\text{re}}=(\text{real}(\lambda_{j})),\lambda_{\text{im}}=(\text{imag}(\lambda_{j}))\in\mathbb{R}^{d}$ be the aggregated vector of real and imaginary parts of the eigenvalues, respectively. We assemble these values into a real-valued block-diagonal matrix:
$\Lambda=\begin{bmatrix}\operatorname{diag}(\lambda_{\text{re}})&-\operatorname{diag}(\lambda_{\text{im}})\\ \operatorname{diag}(\lambda_{\text{im}})&\operatorname{diag}(\lambda_{\text{re}})\end{bmatrix}\in\mathbb{R}^{2d\times 2d}\ .$
At inference time, the latent state vector $z\in\mathbb{R}^{d}$ is first augmented to $\mathbb{R}^{2d}$ by concatenating a zero vector: $\tilde{z}=[z;0]$. The evolution under the Koopman operator is then computed as
$z_{\text{next}}=P_{\text{inv}}\Lambda P\tilde{z}\$.
Finally, we discard the imaginary part and retain only the first $d$ entries to return a real-valued output of the same dimensionality as the input.
This formulation enables us to simulate rich, expressive dynamics through a fully linear evolution in the lifted Koopman space, while maintaining computational tractability and interpretability. Moreover, this decomposition is particularly amenable to distillation: the structure of the transformation ensures that semantic and temporal coherence is preserved during student training, as the model explicitly disentangles dynamics into magnitude and phase components.
Appendix D Additional Experiments and Analysis
D.1 Ablation Studies
Loss Ablation Study.
We conduct an ablation study to evaluate the contribution of each loss component. As shown in Tab. 5, using only $\mathcal{L}_{\text{pred}}$ or any pairwise combination of the Koopman-related losses still enables the model to learn, as indicated by moderate FID and IS scores. However, incorporating all three losses together, including $\mathcal{L}_{\text{Koopman}}$, creates a synergistic effect—reducing the FID from around 10 to 7.83 and achieving a higher IS than any partial combination. Finally, we highlight the importance of incorporating an adversarial setup, which further improves both FID and IS scores by enhancing perceptual quality.
Note, integrating adversarial losses into alternative distillation methods for diffusion models is often non-trivial and presents significant challenges. Many existing approaches lack the architectural flexibility or training stability required to effectively incorporate such objectives, unlike our proposed KDM. Notably, we extended GET with adversarial training, which improved its FID from 6.91 to 6.13. However, KDM still achieves a superior FID of 4.68, demonstrating its effectiveness beyond prior methods.
Table 5: Koopman losses ablation; bold is best and underscore is second-best.
$\mathcal{L}_{\text{lat}}$ $\mathcal{L}_{\text{pred}}$ $\mathcal{L}_{\text{rec}}$ $\mathcal{L}_{\text{lat}}+\mathcal{L}_{\text{pred}}$ $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{pred}}$ $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{lat}}$ $\mathcal{L}_{\text{Koopman}}$ $\mathcal{L}$ GET + Adv
FID 487 11.2 457 9.53 10.7 11.16 7.83 4.97 6.13
IS 1.00 8.23 1.00 8.82 8.26 8.68 8.83 9.22 8.98
Ablation on Loss Weighting and $\lambda_{\text{adv}}$ Sensitivity
We investigate the effect of different weightings for the components of the Koopman loss, namely $\mathcal{L}_{\text{rec}}$, $\mathcal{L}_{\text{lat}}$, and $\mathcal{L}_{\text{pred}}$, as well as the influence of the adversarial weight $\lambda_{\text{adv}}$. The results indicate that the model is robust to moderate changes in these loss weights, with FID scores remaining within a narrow range (4.94–5.03). For the adversarial term, we varied $\lambda_{\text{adv}}\in\{0.5,1.0,0.25\}$ and observed that $\lambda_{\text{adv}}=1.0$ achieved the best trade-off between fidelity and stability, yielding an FID of 4.57 compared to 4.98 for $\lambda_{\text{adv}}=0.5$ and 4.67 for the default $\lambda_{\text{adv}}=0.25$ setting used in the main paper. These findings suggest that KDM remains stable under different loss scalings and that its performance is not overly sensitive to the choice of $\lambda_{\text{adv}}$, demonstrating strong robustness to training hyperparameters.
Empirical Investigation of Koopman Matrix Size.
We evaluate the impact of varying the size of the Koopman matrix. As shown in Tab. 6, smaller matrices tend to generalize better than larger ones. This result is somewhat counterintuitive, as the scaling experiments in Sec. 6.2 demonstrate a clear trend where increasing the number of parameters typically improves performance. We hypothesize that smaller matrices generalize more effectively, and we leave a deeper understanding of this phenomenon—through the lens of dynamical systems theory and empirical analysis—for future research.
Table 6: FID scores across matrix sizes on the CIFAR-10 unconditional task; bold is best.
Size 1024 4096 9216
FID 5.08 7.24 9.41
D.2 Training Stability and Adversarial Loss Convergence
To evaluate the training stability of KDM, we compare the generator and discriminator loss dynamics of KDM and GET across uniformly sampled training steps. Despite the inclusion of an adversarial component in KDM, both losses exhibit stable and convergent behavior throughout training, indicating that the adversarial term does not introduce significant instability in practice. Fig.6 illustrates the generator and discriminator loss trajectories for both KDM and GET. We discard the 0% step due to corrupted values in the raw logs. Overall, the results confirm that KDM maintains stable optimization behavior comparable to GET, while benefiting from enhanced perceptual quality.
010102020303040405050606070708080909010010002244Training Progress (%)Loss ValueGenerator and Discriminator Loss Convergence for KDM and GETKDM Gen. LossKDM Disc. LossGET Gen. LossGET Disc. Loss
Figure 6: Training loss convergence comparison between KDM and GET. Despite including an adversarial loss, KDM maintains stable generator and discriminator convergence comparable to GET.
D.3 Additional Qualitative Results
We include additional qualitative results of our method, extending the evaluation presented in the main text on more datasets and testing scenarios.
Additional Comparison of Student vs Teacher.
In Figs. 7, 8, 9, and 10, we present generations from our model alongside those of the original EDM model. Corresponding blocks indicate samples generated from the same initial noise.
Figure 7: Unconditional Generation of FFHQ. Sampled images using 79 NFEs with EDM karras2022elucidating (left), and using 1 NFE with KDM (right).
Figure 8: Unconditional Generation of FFHQ. Sampled images using 79 NFEs with EDM karras2022elucidating (left), and using 1 NFE with KDM (right).
Figure 9: Unconditional Generation of CIFAR-10. Sampled images using 35 NFEs with EDM karras2022elucidating (left), and using 1 NFE with KDM (right).
Figure 10: Conditional Generation of CIFAR-10. Sampled images using 35 NFEs with EDM karras2022elucidating (left), and using 1 NFE with KDM (right).
Additional Local Structure Analysis.
We extend the experiment from Sec. 6.1 to AFHQv2, conditional CIFAR-10, and unconditional CIFAR-10 datasets, as shown in Fig. 11, Fig. 12, and Fig. 13, respectively. The results reinforce the conclusions of the main experiment: both EDM and our distilled model exhibit local semantic coherence, preserving features such as position, color, identity, and more—even when traversing farther from the original point in noise space.
$\sigma{=}$0 0.05 0.1 0.2 0.3 0.4 0.5 0.6 $\sigma{=}$0 0.05 0.1 0.2 0.3 0.4 0.5 0.6
Refer to caption Refer to caption
Figure 11: Unconditional AFHQv2: Comparison of neighborhood exploration across different noise levels $\sigma$. Left: EDM karras2022elucidating . Right: our method. Each image corresponds to a small perturbation around the original noise sample.
$\sigma{=}$0 0.05 0.1 0.2 0.3 0.4 0.5 0.6 $\sigma{=}$0 0.05 0.1 0.2 0.3 0.4 0.5 0.6
Figure 12: Conditional CIFAR-10: Comparison of neighborhood exploration across different noise levels $\sigma$. Left: EDM karras2022elucidating . Right: our method. Each image corresponds to a small perturbation around the original noise sample.
$\sigma{=}$0 0.05 0.1 0.2 0.3 0.4 0.5 0.6 $\sigma{=}$0 0.05 0.1 0.2 0.3 0.4 0.5 0.6
Refer to caption Refer to caption
Figure 13: Unconditional CIFAR-10: Comparison of neighborhood exploration across different noise levels $\sigma$. Left: EDM karras2022elucidating . Right: our method. Each image corresponds to a small perturbation around the original noise sample.
Appendix E Additional Details
E.1 Training and Sampling Procedures Psudo-codes
The training process of the model follows a Koopman-inspired backward mapping approach, augmented with adversarial supervision. Given a clean-noisy pair $(x_{0},x_{T})$, the model learns a latent Koopman operator that evolves the noisy encoding backward to approximate the clean encoding, while a decoder reconstructs the input from the latent space. The model also supports class-conditional dynamics via Koopman control.
Algorithm 1 Training Procedure
1:Training pairs $(x_{0},x_{T})$, labels $y$ (optional), model $E_{\phi}$, $E_{\psi}$$C_{\eta}$, $D_{\psi}$ discriminator $D_{\gamma}$
2:for each training step do
3:  Encode $z_{0}\leftarrow E_{\phi}(x_{0},0,y)$
4:  Encode $z_{T}\leftarrow E_{\varphi}(x_{T},T,y)$
5:  Add noise: $\tilde{z}_{0}\leftarrow z_{0}+\epsilon_{0}$, $\tilde{z}_{T}\leftarrow z_{T}+\epsilon_{T}$, where $\epsilon_{T},\epsilon_{0}\sim\mathcal{N}(0,0.4)$
6:  Push via Koopman: $z_{0}^{\text{push}}\leftarrow C_{\eta}(\tilde{z}_{T})$
7:  if conditional Koopman then
8:   Add control: $z_{0}^{\text{push}}\leftarrow z_{0}^{\text{push}}+C_{\mu}(y)$
9:  end if
10:  Decode: $\hat{x}_{0}\leftarrow D_{\psi}(\tilde{z}_{0},0,y)$
11:  Decode pushed: $\hat{x}_{0}^{\text{push}}\leftarrow D_{\psi}(z_{0}^{\text{push}},0,y)$
12:  Compute latent loss: $\mathcal{L}_{\text{lat}}\leftarrow\|z_{0}-z_{0}^{\text{push}}\|^{2}$
[…]
14:  Compute full pass loss: $\mathcal{L}_{\text{pred}}\leftarrow d(x_{0},\hat{x}_{0}^{\text{push}})$
15:  Compute model adversarial loss: $\mathcal{L}^{G}_{\text{adv}}\leftarrow CE(1,D_{\gamma}(\hat{x}_{0}^{\text{push}}))$
16:  Compute discriminator loss: $\mathcal{L}^{D}_{\text{adv}}\leftarrow CE(1,D_{\gamma}(x_{0}))+CE(0,D_{\gamma}(\hat{x}_{0}^{\text{push}}))$
17:  model Combine losses:
$\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{pred}}+\mathcal{L}_{\text{lat}}+w_{\text{adv}}\cdot\mathcal{L}^{G}_{\text{adv}}$
18:  Update $E_{\phi}$, $E_{\varphi}$, $C_{\eta}$, $C_{\mu}$, $D_{\psi}$ and $D_{\gamma}$ via backpropagation
19:end for
$d$ is a distance measure such as MSE or LPIPS. We utilize LPIPS. $CE$ is cross-entropy loss. At inference time, the model can sample clean images by reversing from random Gaussian noise $x_{T}$, applying the Koopman operator in latent space, and decoding to the image space.
Algorithm 2 Sampling Procedure
1:Koopman model: $E_{\phi}$, $C_{\eta}$, $C_{\mu}$, $D_{\psi}$, labels $y$ (optional)
2:Sample noise $x_{T}\sim\mathcal{N}(0,\sigma^{2})$
3:Encode: $z_{T}\leftarrow E_{\varphi}(x_{T},T,y)$
4:Push: $z_{0}^{\text{push}}\leftarrow C_{\eta}(z_{T})$
5:if conditional Koopman then
6:  Add control: $z_{0}^{\text{push}}\leftarrow z_{0}^{\text{push}}+C_{\mu}(y)$
7:end if
8:Decode: $\hat{x}_{0}\leftarrow D_{\psi}(z_{0}^{\text{push}},0,y)$
9:return Clean sample $\hat{x}_{0}$
E.2 Datasets Collection
For the Checkerboard dataset, we sampled using 10 NFEs with EDM karras2022elucidating or flow matching lipman2022flow . Following geng2023one , we sampled CIFAR-10 with 35 NFEs from EDM. For FFHQ and AFHQv2, we used 79 NFEs using EDM, as recommended.
E.3 Method Implementation Details
[…]
Checkerboard.
For this dataset, we build upon the implementation of lipman2023flow . Both $x_{0}$ and $x_{T}$ are encoded using the same denoising architecture as described in their work. The decoding of $z_{0}$ is performed using the inverse of this module. The Koopman operator is implemented as a simple linear layer. Complete implementation details are available in the shared code.
CIFAR-10, FFHQ, and AFHQv2.
We adopt the SongUNet architecture from karras2022elucidating for encoding $x_{0}$ and $x_{T}$, as well as for decoding $z_{0}$. To reduce latent dimensionality, the output channels of the network are set to 1, effectively reducing the latent space size by a factor of 3. The Koopman operator is implemented either as a standard linear transformation or using our custom Fast Koopman module. To maintain architectural fairness, we downscale all networks to match the total parameter count of the original Karras network karras2022elucidating and the GET distillation model geng2023one . The FFHQ and AFHQv2 configurations follow a similar encoder-decoder structure with one key difference: a linear projection layer is appended to the end of the encoder and the beginning of the decoder to further control the latent space dimensionality. All other aspects remain consistent across datasets. Note that the discriminator is trained using a separate optimizer, akin to the Koopman module’s optimizer.
Hyperparameters and Training.
We provide complete training scripts along with plug-and-play configuration files in the supplementary material. An overview of the key hyperparameters used for each dataset is summarized in Tab. 7. Overall, our method requires minimal hyperparameter tuning and we did not perform any hyperparameter search.
  • Unet Out Channels: Controls the number of output channels in SongUNet.
  • Unet Model Channels: Determines the internal channel width, affecting the model’s capacity.
  • Noisy Latent Injection: A Gaussian noise term $\epsilon\sim\mathcal{N}(0,1)$ added to latents (as described in the pseudocode) to improve generalization and mitigate numerical instabilities.
  • Adversarial Weight: Applied to the adversarial loss term in our generator loss; not used for the discriminator loss directly.
  • Linear Module: Indicates the presence of an additional linear layer for compressing the latent space, applied at the output of the encoder and the input of the decoder.
Table 7: Model Hyperparameters Across Datasets
Configuration Checkerboard CIFAR-10 FFHQ AFHQv2
Batch Size 4096 128 64 64
Learning Rate (LR) 0.0003 0.0003 0.0003 0.0003
Iterations 800k 800k 800k 800k
Unet Out Channels 1 1 1 1
Unet Model Channels 64 64 32 32
Noisy Latent Injection 0.4 0.4 0.4 0.4
Adversarial Weight 0.01 0.01 0.01 0.01
Linear Module No No Yes Yes
Latent Dim Size 1024 1024 512 512
Table 8: Human Evaluation of Image Realism. The reported ratio reflects the number of times participants selected the model-generated collage as more realistic, out of a total of 300 evaluations.
EDM (Teacher) KDM (Student)
Ratio $45\%$ $55\%$
E.4 Human Evaluation Experiment
While the Fréchet Inception Distance (FID) is widely used to evaluate the quality of images generated by generative models, recent studies have highlighted its limitations in aligning with human judgments of image realism. For instance, Jayasumana et al. jayasumana2024rethinking demonstrate that FID can contradict human evaluations, particularly in assessing perceptual quality and semantic alignment, due to its reliance on Inception-v3 features and the assumption of Gaussian distributions. Similarly, Otani et al. otani2023toward emphasize that automatic metrics like FID often fail to capture the nuanced aspects of human perception, underscoring the need for standardized human evaluation protocols. Furthermore, Kynkäänniemi et al. kynkaanniemi2022role reveal that FID can be manipulated by aligning top-N class histograms without genuinely improving image quality, indicating potential biases in the metric. These findings suggest that while FID provides a useful quantitative signal, it may not fully encapsulate human perceptions of image realism, highlighting the importance of incorporating human evaluations when assessing generative models.
Therefore, in the main paper, we report the results of a human evaluation experiment designed to compare our method against EDM. Specifically, we curated 10 image collages ($\times6$) of human faces generated by the teacher model (EDM, 79 NFEs) and 10 collages generated by our student model (1 NFE), using samples from the FFHQ dataset. Each comparison slide presents one collage from each model, displayed side-by-side as illustrated in Fig. 4, with the order randomized. We then asked 30 human raters to evaluate 10 such slides and choose, for each, which collage appeared more realistic. If both appeared equally realistic, they were instructed to make a random choice. This process yielded 300 total decisions. We report in the main paper and in Tab. 8, the proportion of times each model was preferred. Finally, we attach to the appendix both the slides that we sent to the practitioners and the Excel that aggregates the results. The results indicate that decisions were made almost at random, suggesting that participants had difficulty distinguishing between the student and teacher generations. This implies that, despite existing FID gaps, human evaluators found it hard to perceive meaningful differences. Participants reported that both collages included highly realistic images, as well as images with noticeable artifacts that appeared unrealistic, in their judgment.
E.5 Cost and Efficiency of Offline Distillation
An important consideration in offline distillation is the total computational cost, which includes both the training of the student model and the one-time cost of generating the teacher-produced noise–data pairs. While this pair generation step incurs an initial overhead, it is performed only once and can be reused across multiple student architectures or experiments, effectively amortizing its cost. Following the analysis of (geng2023one, ), which explicitly incorporates the cost of data-pair generation into its complexity assessment, the overall computational load of offline distillation remains favorable compared to online approaches. This is primarily because online distillation requires repeated access to the full teacher model throughout training, resulting in substantially higher neural function evaluations and memory usage. In our experiments on the main paper, we observe that even when accounting for the data generation step, our method achieves a markedly better trade-off between cost and performance than prior offline frameworks such as GET. These results collectively support that, despite the initial preprocessing overhead, offline distillation constitutes an efficient and scalable alternative for practical deployment. To see the full analysis between online and offline approaches, please refer to geng2023one .
E.6 Parameter Scaling
We report the full results of the scaling experiment shown in Sec . 6.2, including IS score from the parameter scaling comparison in Tab. 9. The results demonstrate that our approach consistently outperforms GET in terms of FID across all model sizes, except for IS in 20M, indicating superior perceptual quality. Notably, the improvement is more pronounced as the model size increases, with our 84M model achieving an FID of 4.89 compared to 7.19 for GET. In terms of IS, our method matches or surpasses GET, with particularly strong results at larger scales. These results underscore the scalability and effectiveness of our approach.
Table 9: FID and IS scores for different model sizes, comparing our approach with GET.
20M 38M 84M
Ours GET Ours GET Ours GET
FID 10.2 10.72 5.04 8.00 4.89 7.19
IS 8.50 8.69 9.12 9.03 9.11 9.09
E.7 Extended Outlier Analysis
In Sec. 6.1, we present the results of our model’s component (left side) in the experiment shown in Fig. 14. Outlier detection was performed using the DBSCAN algorithm with a distance threshold of $0.15$. Interestingly, our model exhibits a slightly increased number of outliers. We hypothesize that this may be due to early stopping during training, as our primary objective was not to maximize performance on this particular dataset, but rather to gain insight and intuition into the nature of the learned dynamics. Finally, we observe that the outlier phenomenon is inherited from the teacher model (right side), suggesting that it is a more general characteristic of the generative dynamics rather than an artifact introduced by the distillation process.
[…]
Figure 14: Outlier analysis on generated checkerboard patterns. The two left images are our model (student), and the two images on the right are from the flow matching model (teacher).
Appendix F Additional Discussion of Future Work and Current Limitations
In this study, we took a significant step toward leveraging dynamical systems theory for offline distillation, demonstrating its strong potential in this domain. In this section, we outline several future research directions that we believe are both promising and intellectually compelling.
Standalone Koopman Generative Model.
Koopman theory enables modeling nonlinear dynamical systems in a transformed space where they can be analyzed linearly. This raises an intriguing question: Can the Koopman framework serve as a standalone generative model, capable of learning data distributions from scratch? While the question remains open, exploring this direction could yield a novel and theoretically grounded framework for generative modeling.
Utilization of Full Trajectory Information.
Our current offline setup relies solely on noise-data pairs extracted from pre-trained teacher models. However, the diffusion process inherently includes intermediate steps that are discarded. We hypothesize that leveraging this full trajectory information could enhance trajectory alignment and improve student performance. Moreover, Koopman theory naturally accommodates multi-step signals, making this a promising direction for further investigation. Additionally, the current framework does not support a quality-versus-sampling-steps trade-off, which some online methods do enable. Future work incorporating the full diffusion trajectory could integrate this flexibility into our approach.
Incorporating Broader Koopman Theory Literature.
The connection between Koopman theory and diffusion model distillation opens the door to a rich body of theoretical and practical tools. On the theoretical side, works on sample complexity and error bounds for Koopman operators philipp2023error ; chen2018sample ; muthirayan2021online may provide insights into convergence guarantees. Information-theoretic perspectives lozano2022information could offer new modeling constraints or regularization schemes. On the practical side, methods for rare-event sampling zhang2022koopman may improve long-tail generation capabilities. Finally, while the current mathematical formulation supports discrete Koopman evaluation, the underlying theory naturally extends to continuous settings as well. Exploring these directions could substantially extend and generalize the impact of our framework.
Enhancing the Adversarial Setup.
While adversarial components have been shown to be critical in many distillation methods yin2024improved ; kim2024consistency , our current implementation uses a basic discriminator architecture with standard loss functions and training schemes. We believe that integrating more advanced adversarial paradigms could significantly boost performance, potentially closing the gap with state-of-the-art online distillation methods.
Decision Boundary Insights.
In our synthetic experiments, we identified a common cause underlying failure modes or “outliers” in the generative dynamics. Further investigation is needed to determine whether this phenomenon generalizes to more complex datasets. If it does, it could guide the development of strategies such as uncertainty estimation, selective sampling to avoid generating low-quality outputs, incorporating structural constraints into the latent space to prevent such failures, or introducing auxiliary self-supervised objectives—such as contrastive learning—to enhance the model’s ability to distinguish between semantically similar groups.
Toward Broader Applications.
Due to limited computational resources, we were unable to extend our offline distillation framework to more complex tasks, such as text-to-image generation, that require large-scale compute infrastructures. It will be interesting to explore how our method can scale or adapt to other domains and tasks. Additionally, because our method makes no assumptions about the initial and terminal states of the underlying dynamical system, we believe it is applicable to a broader class of generative models, including DDBMs zhou2023denoising , stochastic interpolants albergo2023stochastic , and others de2021diffusion .